# No Free Lunch from Random Feature Ensembles

### v1

@@ A.2 Gaussian Universality Ansatz and the connection to Linear Random Feature Ridge Regression
Appendix B Proof of Theorems 1 and 2
@@
$\operatorname{Df}_{2}$ $<\operatorname{Df}_{2}$ $\gamma_{2}$ $<\gamma_{1}$ 00 0 $<\rho<1$ (B.1)
Furthermore, the inequality $\operatorname{tf}_{2}(\kappa)\leq\operatorname{tf}_{1}(\kappa)$ holds, with strict inequality when the target has a learnable component.
B.1 “Bigger is Better” Theorems
We begin by proving theorem 1. We note that it suffices to prove that $E_{g}^{K}$ decreases monotonically with $K$, $N$, and $P$ separately, since any transformation $(K,N,P)\to(K^{\prime},N^{\prime},P^{\prime})$ can be taken in steps $K\to K^{\prime}$, $N\to N^{\prime}$, $P\to P^{\prime}$.
@@
$E_{g}^{K}=(1-\frac{1}{K})\operatorname{Bias}_{z}^{2}+\frac{1}{K}E_{g}^{1}$ (B.2)
With $\kappa_{2}$ and $N$ fixed, we see that only the prefactor of $\frac{1}{1-\gamma_{2}}$ in $E_{g}^{1}$ will be affected, so that $E_{g}^{1}$ decreases with $P$. Note also that $\gamma_{1}$ is a decreasing function of $P$. With $\kappa_{2}$ and $N$ fixed, it follows that $E_{g}^{1}$ decreases with $P$. Finally, because $K$ is fixed, $E_{g}^{K}$ will decrease with $P$.
Monotonicity with $N$
Consider a transformation $N\to N^{\prime}$ with $N^{\prime}>N$. Examining eq. 11, we see that it is always possible to increase the ridge $\lambda$ so that $\kappa_{2}$ remains constant. With $\kappa_{2}$, $P$ fixed, $\operatorname{Bias}_{z}^{2}$ is fixed as well. From eq. B.2, it then suffices to show that $E_{g}^{1}$ decreases. To see this, recall that $\rho$ is an increasing function of $N$, and $\gamma_{1}$ is a decreasing function of $\rho$. We then have that $\gamma_{1}$ is a decreasing function of $N$, so that the prefactor of $\frac{1}{1-\gamma_{1}}$ in eq. 9 is decreasing with $N$.
B.2 “No Free Lunch” from Ensembles Theorem
We now prove theorem 2. We first recall the form of the error to be:
$E_{g}^{K}=\operatorname{Bias}_{z}^{2}+\frac{1}{K}\left(E_{g}^{1}-\operatorname{Bias}_{z}^{2}\right)$ (B.3)
We define the variable $\nu\equiv\frac{1}{K}$. By analytical continuation, it suffices to show that test risk decreases as $\nu$ increases. Rewriting the self-consistent equation in terms of $\nu$, we have;
$\kappa_{2}=\frac{\lambda N}{(P-\operatorname{Df}_{1}(\kappa_{2}))(\nu M-\operatorname{Df}_{1}(\kappa_{2}))}$ (B.4)
@@
$\nu(E_{g}^{1}-\operatorname{Bias}_{z}^{2})=$ $\nu\left[\frac{-\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})}{1-\gamma_{1}}-\frac{-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})}{1-\gamma_{2}}\right]$ (B.5)
$+$ $\nu\left[\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right]\sigma_{\epsilon}^{2}$ (B.6)
@@
$\frac{d}{d\nu}\left[\nu\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)\right]<0,$ (B.7)
@@
$\frac{d\rho}{d\nu}$ $=M\frac{(\operatorname{Df}_{1}-\operatorname{Df}_{2})}{(\nu M-\operatorname{Df}_{2})^{2}}$ (B.8)
$\frac{d\gamma_{2}}{d\nu}$ $=-\frac{1}{P}(\operatorname{Df}_{1}-\operatorname{Df}_{2})\frac{d\rho}{d\nu}=-\frac{M}{P}\frac{(\operatorname{Df}_{1}-\operatorname{Df}_{2})^{2}}{(\nu M-\operatorname{Df}_{2})^{2}}$ (B.9)
@@
$\frac{d}{d\nu}\left[\nu\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)\right]$ (B.10)
$=\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)+\frac{\nu\frac{d\gamma_{1}}{d\nu}}{(1-\gamma_{1})^{2}}$ (B.11)
$=\frac{(\operatorname{Df}_{1}-\operatorname{Df}_{2})^{2}}{(1-\gamma_{1})(\nu M-\operatorname{Df}_{2})}\left[\frac{1}{P(1-\gamma_{2})}-\frac{M\nu}{P(1-\gamma_{1})(\nu M-\operatorname{Df}_{2})}\right]$ (B.12)
$<0$ (B.13)
@@
$\frac{d}{d\nu}\left[\nu\left(\frac{-\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})}{1-\gamma_{1}}-\frac{-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})}{1-\gamma_{2}}\right)\right]\leq 0,$ (B.14)
we first note that $-\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})$ can be equivalently written as $-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{2}(\kappa_{2})$. The above derivative can then be broken into two parts:
$-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}\frac{d}{d\nu}\left[\nu\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)\right]+\kappa_{2}\operatorname{tf}_{2}\frac{d}{d\nu}\left[\frac{\nu(1-\rho)}{1-\gamma_{1}}\right]$ (B.15)
We have already shown that the derivative in the first term is negative. Furthermore, $-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}\geq 0$, with strict equality holding when the task has a learnable component. To see that the derivative in the second term is negative, note that $\rho$ is an increasing function of $\nu$. Because $\gamma_{1}$ is a decreasing function of $\rho$, $\gamma_{1}$ is therefore a decreasing function of $\nu$. The denominator $1-\gamma_{1}$ inside the derivative increases with $\nu$. Furthermore, the numerator $\nu(1-\rho)$ can be written as $(\operatorname{Df}_{1}-\operatorname{Df}_{2})\frac{\nu}{\nu M-\operatorname{Df}_{2}}$. With $\kappa_{2}$ fixed (so that $\operatorname{Df}_{1}-\operatorname{Df}_{2}>0$ is fixed), this is a strictly decreasing function of $\nu$. It follows that
$\kappa_{2}\operatorname{tf}_{2}\frac{d}{d\nu}\left[\frac{\nu(1-\rho)}{1-\gamma_{1}}\right]\leq 0,$ (B.16)
@@
Appendix C Derivation of Scaling Laws
In this section, we derive the width-bottlenecked scaling laws given in section 5, using methods described in (Atanasov et al., 2024). We assume that the kernel eigenspectrum decays as $\eta_{t}\sim t^{-\alpha}$ and the power of the target function in the modes decays as $\bar{w}_{t}^{2}\eta_{t}\sim t^{-(1+2\alpha r)}$, and examine the regime where $P\gg N$, so that the width of the ensemble members is the bottleneck to signal recovery. We begin by analyzing the self-consistent equation for $\kappa_{2}$, reproduced here for clarity:
$\kappa_{2}=\frac{\lambda N}{(P-\operatorname{Df}_{1}(\kappa_{2})(N-\operatorname{Df}_{1}(\kappa_{2})))}$ (C.1)
@@ Appendix C Derivation of Scaling Laws
$P\kappa_{2}\approx\frac{\lambda}{1-\frac{1}{N}\operatorname{Df}_{1}(\kappa_{2}))}$ (C.2)
@@
[…]
@@
$\operatorname{Df}_{1}(\kappa_{2})\approx\kappa_{2}^{-1/\alpha}\int_{\kappa_{2}^{1/\alpha}}^{\infty}\frac{du}{1+u^{\alpha}}$ (C.4)
@@
$\kappa_{2}P\approx\frac{\lambda}{1-\frac{\kappa_{2}^{-1/\alpha}}{N}\int_{\kappa_{2}^{1/\alpha}}^{\infty}\frac{du}{1+u^{\alpha}}}$ (C.5)
@@
$PN^{-q}\sim\frac{\lambda}{1-N^{\frac{q}{\alpha}-1}\int_{N^{-q/\alpha}}^{\infty}\frac{du}{1+u^{\alpha}}}$ (C.6)
@@
With this scaling for $\kappa_{2}$, we have $\operatorname{Df}_{1},\operatorname{Df}_{2}\sim N$. It is then clear that $\gamma_{2}\to 0$ for $P\gg N$. Furthermore, because $\rho\in[0,1]$, $\gamma_{1}\to 0$ for $P\gg N$. The prefactors of $\frac{1}{1-\gamma_{2}}$ and $\frac{1}{1-\gamma_{1}}$ can therefore be ignored.
@@
[…]
where $u=t\kappa^{1/\alpha}$ and we have made the substitution $\kappa\sim N^{-\alpha}$. We get two contributions to the integral: when $u$ is near $1/N$, we get a contribution (including the prefactor) which scales as $N^{-\alpha}$. When $u$ is away from $1/N$, the integral contributes a constant factor and we get a contribution that scales as the prefactor $N^{-2\alpha r}$.
@@
[…]
@@
$\operatorname{Bias}_{z}^{2}\sim N^{-2\alpha\min(r,1)}$ (C.9)
$\operatorname{Var}_{z}^{2}\sim N^{-2\alpha\min(r,\frac{1}{2})}$ (C.10)
@@
$\operatorname{Bias}_{z}^{2}$ $\sim M^{-2\ell\alpha\min(r,1)}$ (C.11)
$\frac{1}{K}\operatorname{Var}_{z}$ $\sim M^{-\left(1-\ell+2\alpha\ell\min\left(r,\frac{1}{2}\right)\right)}$ (C.12)
@@
C.1 Sample-Bottlenecked Scaling

### v2

@@ A.2 Gaussian Universality Ansatz and the connection to Linear Random Feature Ridge Regression
Appendix B Notions of Deterministic Equivalent Error
Here, we review the notions of deterministic equivalent error relating the “true” error $\mathcal{E}_{g}^{K}$ of a RFRR ensemble to its deterministic approximation $E_{g}^{K}$. While $\mathcal{E}_{g}^{K}$ is a random quantity, depending on the particular realization of the dataset $\mathcal{D}$ as well as the random-feature weights $\{\bm{v}_{n}^{k}\}$. However, many works have found that the stochastic quantity $\mathcal{E}_{g}^{K}$ concentrates about a deterministic equivalent error $E_{g}^{K}$ which depends on the dataset and random features only through their sizes $P$, NN, and $K$ (see eq. 20). The exact relationship between $\mathcal{E}_{g}^{K}$ and $E_{g}^{K}$, however, varies in the literature. Here we describe two notions of deterministic equivalence, and discuss how and when they apply to the relationship $\mathcal{E}_{g}^{K}\simeq E_{g}^{K}$ employed in this work.
To derive a deterministic equivalent error formula, many works examine an “asymptotic” regime, in which both $N,P\to\infty$ with a fixed ratio ratio $P/N\sim\mathcal{O}(1)$. This permits the use of tools from random matrix theory to analyze the learned predictor. This approach is reviewed in (Atanasov et al., 2024), and we defer the reader to the citations therein. Under the gaussian equivalence assumption which we employ, eq. 20 is a straightforward application of the error formula for linear random feature models (Atanasov et al., 2024, 2022; Adlam & Pennington, 2020; Maloney et al., 2022), with the role of the “data” played by the infinite-dimensional features of the RKHS of the limiting kernel $\bm{H}(\bm{x},\bm{x}^{\prime})$. While asymptotic equivalence holds only in the limit $N,P\to\infty$, the resulting formulas often serve as a good approximation at finite $N$ and $P$
A stronger notion of deterministic equivalence between $\mathcal{E}_{g}^{1}$ and $E_{g}^{1}$ was recently proven by Defilippis et al. without the assumption of Gaussian Equivalence, resolving the conjecture of (Simon et al., 2023) for single models ($K=1$). Specifically, Defilippis et al. proved a “dimension-free” deterministic equivalence which guarantees that
$|\mathcal{E}_{g}^{1}-E_{g}^{1}|\leq C\,\bigl(\tfrac{1}{\sqrt{N}}+\tfrac{1}{\sqrt{P}}\bigr)E_{g}\,$,
with high probability, for a factor $C$ which is logarithmically small in $N$ and $P$. Unlike asymptotic deterministic equivalence, dimension-free deterministic equivalence applies at fixed, finite $P$ and $N$. While at the time of writing the dimension-free deterministic equivalence has not been proven the the case of ensembled RFRR ($K>1$), we anticipate that the arguments of (Defilippis et al., 2024) could be extended to the ensembled case.
Appendix C Proof of Theorems 4.1 and 5.1
@@ Appendix B Notions of Deterministic Equivalent Error
$\operatorname{Df}_{2}$ $<\operatorname{Df}_{1}$ $\gamma_{2}$ $<\gamma_{1}$ $0$ $<\rho<1$ (C.1)
We verify all deterministic equivalent error formulas used here by comparing to numerical experiments.
Furthermore, the inequality $\operatorname{tf}_{2}(\kappa)\leq\operatorname{tf}_{1}(\kappa)$ holds, with strict inequality when the target has a learnable component. In each of the proofs that follow, we aim to show that the error of a random-feature ensemble decreases monotonically under certain transformations of the parameters $(K,N,P)\to(K^{\prime},N^{\prime},P^{\prime})$ when the value of the ridge parameter $\lambda$ is set to it’s optimal value. We will argue using the invariance-based method introduced by (Simon et al., 2023). Denote the errors before and after the transformation by $E_{g}^{K}(N,P,\lambda)$ and $E_{g}^{K^{\prime}}(N^{\prime},P^{\prime},\lambda^{\prime})$. We allow here for the regularization $\lambda$ before the transformation to take an arbitrary value. Then, we will prove that there always exists a value $\lambda^{\prime}$ such that $E_{g}^{K^{\prime}}(N^{\prime},P^{\prime},\lambda^{\prime})\leq E_{g}^{K}(N,P,\lambda)$. Proof follows by setting $\lambda$ equal to it’s optimal value $\lambda^{*}$ and noting that $\min_{\lambda}E_{g}^{K^{\prime}}(N^{\prime},P^{\prime},\lambda)\leq E_{g}^{K^{\prime}}(N^{\prime},P^{\prime},\lambda^{\prime})\leq E_{g}^{K}(N,P,\lambda^{*})$. Below, we will prove the existence of $\lambda^{\prime}$ under the desired transformations.
C.1 “Bigger is Better” Theorems
We begin with the setting of the 4.1, so that we must prove error decreases under any transformation $(K,N,P)\to(K^{\prime},N^{\prime},P^{\prime})$ such that $K\leq K^{\prime}$, $N\leq N^{\prime}$, and $P\leq P^{\prime}$. We note that it suffices to prove that $E_{g}^{K}$ decreases monotonically with $K$, NN, and $P$ separately, since any transformation $(K,N,P)\to(K^{\prime},N^{\prime},P^{\prime})$ can be taken in steps $K\to K^{\prime}$, $N\to N^{\prime}$, $P\to P^{\prime}$.
@@
$E_{g}^{K}=(1-\frac{1}{K})\operatorname{Bias}_{z}^{2}+\frac{1}{K}E_{g}^{1}$ (C.2)
With $\kappa_{2}$ and $N$ fixed, we see that only the pre-factor of $\frac{1}{1-\gamma_{2}}$ in $E_{g}^{1}$ will be affected, so that $E_{g}^{1}$ decreases with $P$. Note also that $\gamma_{1}$ is a decreasing function of $P$. With $\kappa_{2}$ and $N$ fixed, it follows that $E_{g}^{1}$ decreases with $P$. Finally, because $K$ is fixed, $E_{g}^{K}$ will decrease with $P$.
Monotonicity with NN
Consider a transformation $N\to N^{\prime}$ with $N^{\prime}>N$. Examining eq. 14, we see that it is always possible to increase the ridge $\lambda$ so that $\kappa_{2}$ remains constant. With $\kappa_{2}$, $P$ fixed, $\operatorname{Bias}_{z}^{2}$ is fixed as well. From eq. C.2, it then suffices to show that $E_{g}^{1}$ decreases. To see this, recall that $\rho$ is an increasing function of $N$, and $\gamma_{1}$ is a decreasing function of $\rho$. We then have that $\gamma_{1}$ is a decreasing function of $N$, so that the pre-factor of $\frac{1}{1-\gamma_{1}}$ in eq. 12 is decreasing with $N$.
C.2 “No Free Lunch” from Ensembles Theorem
We now prove theorem 5.1. We first recall the form of the error to be:
$E_{g}^{K}=\operatorname{Bias}_{z}^{2}+\frac{1}{K}\left(E_{g}^{1}-\operatorname{Bias}_{z}^{2}\right)$ (C.3)
We define the variable $\nu\equiv\frac{1}{K}$. By analytical continuation, it suffices to show that test risk decreases as $\nu$ increases when $M=KN$ is held fixed. Rewriting the self-consistent equation in terms of $\nu$, we have;
$\kappa_{2}=\frac{\lambda N}{(P-\operatorname{Df}_{1}(\kappa_{2}))(\nu M-\operatorname{Df}_{1}(\kappa_{2}))}$ (C.4)
@@
$\nu(E_{g}^{1}-\operatorname{Bias}_{z}^{2})=$ $\nu\left[\frac{-\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})}{1-\gamma_{1}}-\frac{-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})}{1-\gamma_{2}}\right]$ (C.5)
$+$ $\nu\left[\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right]\sigma_{\epsilon}^{2}$ (C.6)
@@
$\frac{d}{d\nu}\left[\nu\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)\right]<0,$ (C.7)
@@
$\frac{d\rho}{d\nu}$ $=M\frac{(\operatorname{Df}_{1}-\operatorname{Df}_{2})}{(\nu M-\operatorname{Df}_{2})^{2}}$ (C.8)
$\frac{d\gamma_{2}}{d\nu}$ $=-\frac{1}{P}(\operatorname{Df}_{1}-\operatorname{Df}_{2})\frac{d\rho}{d\nu}=-\frac{M}{P}\frac{(\operatorname{Df}_{1}-\operatorname{Df}_{2})^{2}}{(\nu M-\operatorname{Df}_{2})^{2}}$ (C.9)
@@
$\frac{d}{d\nu}\left[\nu\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)\right]$ (C.10)
$=\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)+\frac{\nu\frac{d\gamma_{1}}{d\nu}}{(1-\gamma_{1})^{2}}$ (C.11)
$=\frac{(\operatorname{Df}_{1}-\operatorname{Df}_{2})^{2}}{(1-\gamma_{1})(\nu M-\operatorname{Df}_{2})}\left[\frac{1}{P(1-\gamma_{2})}-\frac{M\nu}{P(1-\gamma_{1})(\nu M-\operatorname{Df}_{2})}\right]$ (C.12)
$<0$ (C.13)
@@
$\frac{d}{d\nu}\left[\nu\left(\frac{-\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})}{1-\gamma_{1}}-\frac{-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})}{1-\gamma_{2}}\right)\right]\leq 0,$ (C.14)
we first note that $\rho\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})$ can be equivalently written as $\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})+(1-\rho)\kappa_{2}\operatorname{tf}_{2}(\kappa_{2})$. The above derivative can then be broken into two parts:
$-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}\frac{d}{d\nu}\left[\nu\left(\frac{1}{1-\gamma_{1}}-\frac{1}{1-\gamma_{2}}\right)\right]+\kappa_{2}\operatorname{tf}_{2}\frac{d}{d\nu}\left[\frac{\nu(1-\rho)}{1-\gamma_{1}}\right]$ (C.15)
We have already shown that the derivative in the first term is negative. Furthermore, $\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}\geq 0$, with strict equality holding when the task has a learnable component. To see that the derivative in the second term is negative, note that $\rho$ is an increasing function of $\nu$. Because $\gamma_{1}$ is a decreasing function of $\rho$, $\gamma_{1}$ is therefore a decreasing function of $\nu$. The denominator $1-\gamma_{1}$ inside the derivative increases with $\nu$. Furthermore, the numerator $\nu(1-\rho)$ can be written as $(\operatorname{Df}_{1}-\operatorname{Df}_{2})\frac{\nu}{\nu M-\operatorname{Df}_{2}}$. With $\kappa_{2}$ fixed (so that $\operatorname{Df}_{1}-\operatorname{Df}_{2}>0$ is fixed), this is a strictly decreasing function of $\nu$. It follows that
$\kappa_{2}\operatorname{tf}_{2}\frac{d}{d\nu}\left[\frac{\nu(1-\rho)}{1-\gamma_{1}}\right]\leq 0,$ (C.16)
@@
Appendix D Derivation of Scaling Laws
In this section, we derive the width-bottlenecked scaling laws given in section 6.2, using methods described in (Atanasov et al., 2024). We assume that the kernel eigenspectrum decays as $\eta_{t}\sim t^{-\alpha}$ and the power of the target function in the modes decays as $\bar{w}_{t}^{2}\eta_{t}\sim t^{-(1+2\alpha r)}$, and examine the regime where $P\gg N$, so that the width of the ensemble members is the bottleneck to signal recovery. We begin by analyzing the self-consistent equation for $\kappa_{2}$, reproduced here for clarity:
$\kappa_{2}=\frac{\lambda N}{(P-\operatorname{Df}_{1}(\kappa_{2})(N-\operatorname{Df}_{1}(\kappa_{2})))}$ (D.1)
@@ Appendix D Derivation of Scaling Laws
$P\kappa_{2}\approx\frac{\lambda}{1-\frac{1}{N}\operatorname{Df}_{1}(\kappa_{2}))}$ (D.2)
@@
$\operatorname{Df}_{1}(\kappa_{2})\approx\int_{1}^{\infty}\frac{t^{-\alpha}}{t^{-\alpha}+\kappa_{2}}dt$ (D.3)
@@
$\operatorname{Df}_{1}(\kappa_{2})\approx\kappa_{2}^{-1/\alpha}\int_{\kappa_{2}^{1/\alpha}}^{\infty}\frac{du}{1+u^{\alpha}}$ (D.4)
@@
$\kappa_{2}P\approx\frac{\lambda}{1-\frac{\kappa_{2}^{-1/\alpha}}{N}\int_{\kappa_{2}^{1/\alpha}}^{\infty}\frac{du}{1+u^{\alpha}}}$ (D.5)
@@
$PN^{-q}\sim\frac{\lambda}{1-N^{\frac{q}{\alpha}-1}\int_{N^{-q/\alpha}}^{\infty}\frac{du}{1+u^{\alpha}}}$ (D.6)
@@
With this scaling for $\kappa_{2}$, we have $\operatorname{Df}_{1},\operatorname{Df}_{2}\sim N$. It is then clear that $\gamma_{2}\to 0$ for $P\gg N$. Furthermore, because $\rho\in[0,1]$, $\gamma_{1}\to 0$ for $P\gg N$. The pre-factors of $\frac{1}{1-\gamma_{2}}$ and $\frac{1}{1-\gamma_{1}}$ can therefore be ignored.
@@
$\kappa_{2}\operatorname{tf}_{1}(\kappa_{2})\sim\int_{1}^{\infty}\frac{t^{-(1+2\alpha r)}}{1+t^{-\alpha}/\kappa_{2}}dt\sim N^{-2\alpha r}\int_{1/N}^{\infty}\frac{u^{-(1+2\alpha r)}}{1+u^{-\alpha}}du$ (D.7)
where $u=t\kappa^{1/\alpha}$ and we have made the substitution $\kappa\sim N^{-\alpha}$. We get two contributions to the integral: when $u$ is near $1/N$, we get a contribution (including the pre-factor) which scales as $N^{-\alpha}$. When $u$ is away from $1/N$, the integral contributes a constant factor and we get a contribution that scales as the pre-factor $N^{-2\alpha r}$.
@@
$-\kappa_{2}^{2}\operatorname{tf}_{1}^{\prime}(\kappa_{2})\sim\int_{1}^{\infty}\frac{t^{-(1+2\alpha r)}}{(1+t^{-\alpha}/\kappa_{2})^{2}}dt\sim N^{-2\alpha r}\int_{1/N}^{\infty}\frac{u^{-(1+2\alpha r)}}{(1+u^{-\alpha})^{2}}du$ (D.8)
@@
$\operatorname{Bias}_{z}^{2}\sim N^{-2\alpha\min(r,1)}$ (D.9)
$\operatorname{Var}_{z}\sim N^{-2\alpha\min(r,\frac{1}{2})}$ (D.10)
@@
$\operatorname{Bias}_{z}^{2}$ $\sim M^{-2\ell\alpha\min(r,1)}$ (D.11)
$\frac{1}{K}\operatorname{Var}_{z}$ $\sim M^{-\left(1-\ell+2\alpha\ell\min\left(r,\frac{1}{2}\right)\right)}$ (D.12)
@@
D.1 Sample-Bottlenecked Scaling